Josie is 11 years older than Macy. What is an equation that relates the age of Josie J and the age of Macy M?
I tried:
J = M + 11

Answer by josgarithmetic(39617)

j for Jose and m for Macy.

j = 11 more than m, "11 years older than Macy".
j=11+m or j=m+11

Either way, your equation is the same (just using slightly different variable names).